### v1

# Learned Free-Energy Functionals from Pair-Correlation Matching for Dynamical Density Functional Theory

Abstract
Classical density functional theory (cDFT) and dynamical density functional theory (DDFT) are modern statistical mechanical theories for modelling many-body colloidal systems at the one-body density level. The theories hinge on knowing the excess free-energy accurately, which is however not feasible for most practical applications. Dijkman et al. [Phys. Rev. Lett. 134, 056103 (2025)] recently showed how a neural excess free-energy functional for cDFT can be learned from inexpensive bulk simulations via pair-correlation matching. In this Letter, we demonstrate how this same functional can be applied to DDFT, without any retraining, to simulate non-equilibrium overdamped dynamics of inhomogeneous densities. We evaluate this on a 3D Lennard-Jones system with planar geometry under various complex external potentials and observe good agreement of the dynamical densities with those from expensive Brownian dynamic simulations, up to the limit of the adiabatic approximation. We further develop and apply an extension of DDFT to a grand-canonical system modeled after breakthrough gas adsorption studies, finding similarly good agreement. Our results demonstrate a practical route for leveraging learned free-energy functionals in DDFT, paving the way for accurate and efficient modeling of non-equilibrium systems.
Introduction—Studying the collective behavior of many-body systems emerging from inter-particle interactions is fundamental in many diverse fields of science, ranging from gas adsorption studies in porous media [1] to semiconductors and protein adsorption [2]. Classical density functional theory (cDFT) [3] is a promising theory for accurately and efficiently modelling such many-body systems at the level of the one-body density profile $\rho(\mathbf{r})$ in thermodynamic equilibrium. cDFT relies on the existence of an excess free-energy functional $\mathcal{F}_{\text{exc}}[\rho(\mathbf{r})]$ which describes the inter-particle interactions, accounting for the non-ideal contribution to the total Helmholtz intrinsic free-energy functional $\mathcal{F}[\rho(\mathbf{r})]$. This functional is generally unknown and various approximations have been constructed ranging from the mean-field approximation [4] for soft interactions to fundamental measure theory [5] for hard-sphere interactions. More recently, neural functionals learned from Monte Carlo simulations have been proposed [6, 7, 8, 9, 10] and exhibit remarkable accuracy.
Dynamical density functional theory (DDFT) [3, 11, 12, 13] is an extension of cDFT to describe non-equilibrium systems as a Fokker-Planck equation on the one-body density profile. DDFT has been applied in diverse systems ranging from simple colloidal fluids [11] to plasmas [14] and microswimmers [15]. A recent review of DDFT applications can be found in [16]. Like cDFT, DDFT also relies on the excess free-energy functional and typically the same equilibrium functional from cDFT is adopted for DDFT under the adiabatic approximation.
In this preprint, we adopt the neural functional proposed by Dijkman et al. [10], that was trained by matching the pair-correlation function from bulk simulations, and evaluate it for DDFT. Despite only having seen bulk densities during training, we show that this functional produces accurate time evolution of $1$D inhomogeneous densities on a $3$D Lennard-Jones system with planar geometry. We further extend DDFT with a particle source term to model grand-canonical systems with non-conserved dynamics, inspired by gas adsorption studies, and we observe similarly good agreement with a hybrid Brownian dynamic simulation.
Pair-correlation matching—Pair-correlation matching [10] is a method for learning an accurate neural excess free-energy functional $\mathcal{F}_{\theta,\text{exc}}$ (where $\theta$ represents neural network parameters) by training exclusively on a dataset of radial distribution functions, rather than from expensive heterogeneous density profiles. The method uses the relationship between the second functional derivative of the excess free-energy functional $\frac{\delta^{2}\mathcal{F}_{\text{exc}}}{\delta\rho(z_{i})\delta\rho(z_{j})}$ and the pair-correlation function $c^{(2)}(\mathbf{r})$, which is related to the radial distribution function $g(r)$ via the Ornstein-Zernike equation ([4], Ch.3), for training. Specifically, a neural network $\mathcal{F}_{\theta,\text{exc}}$ is trained by optimizing a physics-based regularizer [17], that matches its Hessian $\frac{\partial^{2}\mathcal{F}_{\text{exc}}}{\partial\rho_{i}\partial\rho_{j}}$ with the direct correlation function obtained from Monte Carlo simulations of homogeneous bulk systems. This functional is then used to obtain accurate estimates for inhomogeneous equilibrium density profiles under various external fields without having used any inhomogeneous densities during training.
Figure 1: An overview of the proposed neural DDFT framework, and the pair-correlation matching [10] approach. (Bottom) A neural free-energy functional is trained via pair-correlation matching where its second functional derivative is optimized against the direct pair correlation function computed from simulated uniform bulk densities. (Top) The neural excess free-energy functional $\mathcal{F}_{\theta,\text{exc}}$ is used for DDFT to time evolve the inhomogeneous particle density $\rho(\mathbf{r})$ of a 3D Lennard-Jones system with planar geometry via automatic differentiation and numerical integration, towards equilibrium.
Classical dynamical density functional theory—DDFT exists in deterministic and stochastic variants, with the deterministic variant first introduced phenomenologically by Evans [3], and later microscopically derived from the Langevin equations by Marconi and Tarazona  [11], and from Smoluchowski equations by Archer and Evans [12]. An extensive review of DDFT fundamentals can be found in [13]. Deterministic DDFT has the form of a continuity equation where $\partial_{t}\rho(\mathbf{r},t)$ is proportional to the divergence of a vector field as follows,
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}=\Gamma\nabla\cdot\left[\rho(\mathbf{r},t)\nabla\left(\frac{\delta\mathcal{F}[\rho]}{\delta\rho(\mathbf{r},t)}\right)\right],$ (1)
where $\Gamma$ is the mobility coefficient and $\mathcal{F}[\rho]$ is the intrinsic Helmholtz free-energy functional which can be written as $\mathcal{F}[\rho]=\frac{1}{\beta}\int d^{3}\mathbf{r}\rho(\mathbf{r})(\ln(\Lambda^{3}\rho(\mathbf{r}))-1)+\int d^{3}\mathbf{r}\rho(\mathbf{r})V_{\text{ext}}(\mathbf{r})+\mathcal{F}_{\text{exc}}([\rho])$.
In the ideal gas limit, where there are no inter-particle interactions i.e. $\mathcal{F}_{\text{exc}}[\rho]=0$, this equation reduces to a drift-diffusion equation where particles only diffuse and respond to an external potential. The DDFT equation in this standard form, with no velocity or inertial terms, describes purely overdamped dynamics where frictional forces dominate. It also relies on the adiabatic approximation which assumes that the non-equilibrium pair correlation function is equivalent to its equilibrium counterpart at the same one-body density. However, for systems where superadiabatic forces such as memory effects are significant, more advanced frameworks such as power functional theory [18] and superadiabatic DDFT [19] are required. It is also worth noting that since standard DDFT has the form of a continuity equation and conserves the particle number (neglecting the boundaries), it is only a canonical theory in this regard. In practice however it is common to still adopt the same grand-canonical free energy functional from cDFT under the adiabatic approximation. We follow this practice and adopt the equilibrium neural excess-free energy functional from Dijkman et al. [10].
Extension to open systems—Motivated by applications in gas adsorption studies where the adsorbent system is typically connected to a grand-canonical particle bath in a hybrid grand-canonical molecular dynamics approach [20, 21], we develop an analog of DDFT for these hybrid simulations. Specifically, we extend the DDFT framework to open systems where particles can be added or removed via a particle bath in the following form,
Figure 2: Time evolution of the one-body density profile $\rho(z,\tau)$ for a system of interacting $3$D LJ particles with planar geometry subjected to a repulsive Gaussian potential $V_{\text{ext}}(z)$ (grey curve, right axis). The plots compare two neural excess free-energy functionals and an analytical FMT functional, with a reference Brownian dynamics simulation. Initially ($t=0\tau$), the system is at equilibrium with a flat density profile. Upon application of the external potential, the density profile evolves non-trivially, exhibiting a depletion near the potential maximum and eventually approaching a new equilibrium. The neural DDFT approaches (red and blue) exhibit good agreement with Brownian dynamics throughout the temporal evolution, notably with the pair-correlation matching approach (red) not having seen any inhomogeneous densities during training.
$\begin{split}\frac{\partial\rho(\mathbf{r},t)}{\partial t}=\,&\underbrace{\Gamma\nabla\cdot\Bigg{[}\rho(\mathbf{r},t)\nabla\Bigg{(}\frac{\delta\mathcal{F}[\rho]}{\delta\rho(\mathbf{r},t)}\Bigg{)}\Bigg{]}}_{\text{continuity equation}}\\ &+\underbrace{(1-\lambda(\mathbf{r}))\Gamma^{\prime}\Bigg{(}\mu-\frac{\delta\mathcal{F}[\rho]}{\delta\rho(\mathbf{r},t)}\Bigg{)}\rho(\mathbf{r},t)}_{\text{growth term}},\end{split}$ (2)
where $\lambda(\mathbf{r})$ is a masking function such that $\lambda(\mathbf{r})=0$ for $\mathbf{r}=\mathbf{r}_{i}$, the locations where a particle bath is to be connected, and is equal to 1 elsewhere. $\Gamma^{\prime}$ is the corresponding mobility coefficient that controls the rate of the non-conserved dynamics. This model is similar to that of Thiele et al. [22], where they used it to model the evaporation of a colloidal suspension into a vapor reservoir. We instead interpret our model using a Wasserstein-Fisher-Rao gradient flow perspective which we expand on in the End Matter.
Figure 3: The final equilibrium density profiles and their RMSE over time evolution under complex external potentials (see SM [23] for the movie version). (a)-(d) A comparison of the final 1D equilibrium density profiles for a three-dimensional LJ system with planar geometry under different external potentials constructed from sums of repulsive Gaussian and steep well potentials. (e)-(h) show the corresponding RMSE deviation from the Brownian dynamics results over time evolution. The densities obtained with the neural functionals are significantly more accurate than those with the analytical functional, notably with the pair-correlation matching approach (red) not having seen any inhomogeneous densities during training.
Implementation—We now describe the numerical implementation of our neural DDFT approach and compare with a Brownian dynamics (BD) reference simulation (see SM [23], Sec.3), where BD is chosen rather than (ballistic) molecular dynamics to be consistent with the overdamped nature of DDFT. We numerically integrate the DDFT equation (eq. 1), using the neural pair-correlation functional, to obtain time-dependent density profiles of $3$D Lennard-Jones (LJ) systems with planar geometry under an external potential $V_{\text{ext}}$ (Fig. 1). We define a $1$D density grid of $10\sigma$ length with a spacing of $\frac{\sigma}{100}$ for the solver, with periodic boundaries. We initialize the density grid with the same uniform initial density as the reference BD simulation and apply the same external potential. We simulate the PDE with a time-step of $10^{-4}\tau$ until $t=4\tau$, where $\tau=\sqrt{\frac{m\sigma^{2}}{\epsilon}}$ in reduced LJ units. The gradients are computed using finite-differences and time integrated using an Euler scheme, as implemented in the numerical Py-PDE Python package [24]. A psuedospectral solver can also be used for scenarios that require more numerical precision [25]. The mobility coefficient $\Gamma$ is chosen to be the same as in BD. For computing $\frac{\delta\mathcal{F}_{\text{exc}}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}$, we use automatic differentiation over the excess neural free-energy functional.
Experiments—We run two sets of “experiments” to evaluate the accuracy and efficiency of neural DDFT from pair-correlation matching with respect to BD, one on a closed system with planar geometry and conserved number of particles under various external potentials, and the other on a system with a source and sink with varying number of particles, inspired by a gas adsorption process.
Closed system: We use a 3D system of LJ particles with planar geometry, under external potentials that are mixtures of repulsive Gaussian and steep well potentials as in [10]. In addition to the neural functional from pair-correlation matching (denoted by $\mathcal{F}_{\theta,\text{exc}}^{(2)}$), we also train a functional using single-body direct correlation matching (denoted by $\mathcal{F}_{\theta,\text{exc}}^{(1)}$) (see SM [23], Sec.1A). Both functionals use the same convolutional neural network with periodic and dilated convolutions, each with a kernel size of 3, a dilation of 2, and 6 layers. The layer dimensions are set to $N_{\text{channels}}$ = [16, 16, 32, 32, 64, 64], with average pooling with kernel size 2 applied after each layer. The models are trained for 5000 epochs, each taking less than an hour on an RTX 4070 GPU for convergence.
For training $\mathcal{F}_{\theta,\text{exc}}^{(2)}$, we simulated $1000$ uniform densities and measured their radial distribution functions with $\frac{\sigma}{100}$ grid-size and chemical potentials in the range -4 to 0.5. $\mathcal{F}_{\theta,\text{exc}}^{(1)}$ is trained on $800$ non-uniform densities [10] with $\frac{\sigma}{100}$ grid-size and chemical potentials in the range -4 to 0.5. We also evaluate against an analytical functional $\mathcal{F}^{\text{MF}}_{\text{exc}}$ which models the attractive contribution using mean-field approximation and the hard-sphere contribution using White-Bear II version of FMT [26], as implemented in PyDFTlj [1].
Our experiments show that the neural free energy functionals $\mathcal{F}_{\theta,\text{exc}}^{(1)}$ and $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ are both significantly better than the $\mathcal{F}^{\text{MF}}$ functional in accuracy across time, as shown in. Fig. 2 & Fig. 3. The convergence of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is similar to $\mathcal{F}_{\theta,\text{exc}}^{(1)}$ despite never having seen inhomogenous densities during training, except in Fig. 3(d) where the density near the wells is pushed above the maximum bulk value of $0.67\sigma^{-3}$ seen during training of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$. The errors for all the DDFT approaches for time intervals $t<1\tau$ are high which we attribute to overly fast dynamics arising from the adiabatic approximation [19]. More notably, the dynamical density profiles from DDFT are several orders of magnitude faster to compute with DDFT than with BD simulation: while running $5000$ trials of BD takes about a day on a single-core CPU, the DDFT approach on the same computer takes less than 6 minutes (see SM [23], Sec.5 for runtime comparisions).
Open system: To evaluate the pair-correlation matching functional on an open system, we construct a system modeled after “breakthrough” experiments used in gas adsorption studies [27, 28]. In these studies, typically a column filled with adsorbent material (e.g. nanoporous metal-organic framework (MOF) is exposed to a gas stream (e.g. CO2), where the breakthrough of the gas is detected at the outlet to estimate the adsorption capacity and kinetics of the material. We adopt such a setup by extending the closed Lennard-Jones system from the previous experiment with two particle baths, one at a high $\mu=0.5$ in $z\in[0,2]\sigma$ to simulate an LJ particle source and the other at extremely low $\mu=-100$ in $z\in[8,10]\sigma$ to simulate an LJ particle sink. The difference in $\mu$ causes particles to flow from the source to the sink through the region $z\in[2,8]\sigma$, in which we apply, in addition, a planar potential to mimic two MOF planes from the IRMOF series [29] (see SM [23], Sec.4).
Figure 4: Breakthrough simulations using neural DDFT and Browninan dynamics (see SM [23] for the movie version). (a) LJ particles from the source move towards the sink under two different MOF-like repulsive external potentials. The dynamics of the density evolution and the final evolved density (visualized) obtained from DDFT are in close agreement with BD. (b) Time evolution of the output to input flux ratio, showing the approach towards steady-state or breakthrough condition. The particles reach breakthrough earlier under the MOF-2 external potential. The curves obtained from DDFT are in close agreement with BD, while being smoother and taking only $2\%$ of the compute time of BD.
We simulate this system using our open DDFT equation (eq. 2) for a duration of $150\tau$, with a timestep of $10^{-3}\tau$ and grid spacing of $\frac{\sigma}{32}$. To obtain a breakthrough curve we compute the time-dependence of the ratio $J(z_{2},t)/J(z_{1},T)$ of the particle fluxes $J(z_{2},t)$ close to the output at $z_{2}=7.5\sigma$ and $J(z_{1},T)$ close to the input at $z_{1}=2.5\sigma$ at late time $T=145\tau$ where steady-state is (almost) reached. The particle flux from DDFT is computed as $J(z,t)=-\Gamma\rho(z,t)\nabla\frac{\delta\mathcal{F}[\rho(z)]}{\delta\rho(z)}$, for which we also apply a Savitsky-Golay filter [30] over the spatial gradient to smooth out noise arising from numerical errors. We compare this with a BD simulation where the particle baths are modelled using hybrid Grand-Canonical Monte Carlo moves (see SM [23], Sec.3). The BD flux is computed as the net number of particles crossing the planes at $z_{1}$ and $z_{2}$ per unit time per unit area, averaged over $500$ BD runs. As observed in Fig. 4, the dynamics of both the BD and DDFT simulations agree closely and reach breakthrough while the DDFT simulation takes only $2\%$ of the compute time of the BD simulation (see SM [23], Sec.5 for runtime comparisions). The potential from MOF-2 (red), which corresponds to a MOF with larger pore size, presents a lower particle barrier which gives rise to a faster evolution to the steady state. Note that since the GCMC moves are run in between each BD update while the open DDFT equation updates are simultaneous, we do not expect exact dynamical agreement though this can be alleviated to some extent with the choice of $\Gamma^{\prime}$ in the growth term.
Conclusion—To summarize, we have presented results that the neural excess free-energy functional trained via pair-correlation matching can be directly applied to DDFT without retraining, with high accuracy relative to an analytical free-energy approximation and with high efficiency relative to BD simulation for conserved dynamics, as well as for non-conserved dynamics via a novel gradient flow based extension inspired by gas adsorption simulations. Future work could focus on the extension of this method to systems with full $3$D geometry and anisotropic particles, and extending it to incorporate superadiabatic forces [31], towards practical applications in modelling non-equilibrium systems.
Acknowledgements—We would like to thank David Dubbeldam for useful discussions.
Data availability—Code and data to reproduce all results will be made available upon publication.